OpenStudy (anonymous):

Determine whether (0, -3) is a solution of 7x - 2y = 7

jimthompson5910 (jim_thompson5910):

7x - 2y = 7 7(0) - 2(-3) = 7 0 + 6= 7 6=7 ... which is false So (0,-3) is NOT a solution.

OpenStudy (anonymous):

7*0 - 2*(-3) = 6 which is not 7